# FedSEA-LLaMA: A Secure, Efficient and Adaptive Federated Splitting Framework for Large Language Models

### v2

@@ Introduction
To investigate the impact of split point selection and noise scale on the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5. Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FedSEA-LLaMA is shown in Table  3. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero. These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
@@
Figure 5(b) shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second. The context length is 500. It can be observed that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving a speed more than seven times faster. Without the KV cache, however, the inference speed decreases significantly as the query length increases, primarily due to the continuously growing number of K and V vectors. In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing previously generated K and V vectors. This effectively avoids redundant computations. When the query length reaches 4000, the inference speed of FedSEA-LLaMA with KV cache is 8.2 times that without it. As query length increases further, this advantage continues to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
@@
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources.
@@
Table 4 shows F1a/EM scores across different partitioning schemes. Performance variations across partitions are minimal, indicating that FedSEA-LLaMA is robust to different split configurations. This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
@@ Conclusion
In this work, we proposed FedSEA-LLaMA, a novel federated split learning framework tailored for large language models. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more base models and deploy FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
@@ Appendix A Noise Added to Hidden States
Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We use LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2. Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2.
@@ Appendix C Parallel Training
Strategy Sequential training    Client-batch training    Server-hierachical training      Centralized
num. of clients 2 2 4 8 2 4 8 -
@@
Appendix D Parameter Setting
@@
Ablation Study on Attention Mask

### v3

@@ Introduction
To assess how split points and noise scale affect the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth ($Blocks_{A}$ = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). Attacks use full-parameter fine-tuning (lr=2e-5), with reconstruction quality measured by BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FedSEA-LLaMA are shown in Table  3. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLM blocks in addition to the embedding layer, all attack metrics approach zero. These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
@@
Figure 5(b) shows a comparison of the time consumption between transferring the attention-mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second(context length=500). We can see that transferring the compressed information of the attention-mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving over 7× speedup. Without the KV cache, the inference speed decreases significantly with longer queries, due to the increasing number of K and V vectors. In contrast, with the KV cache technique, the client only needs to send the intermediate vector of the current single token to the server, while reusing previously generated K and V vectors. This effectively avoids redundant computations. When the query length reaches 4000, the inference speed of FedSEA-LLaMA with KV cache is 8.2 times that without it. As query length increases further, this advantage continues to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
@@
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ ($Blocks_{A}$ = 1, 2, 3 and $Blocks_{C}$ = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources. Table 4 shows F1a/EM scores across different partitioning schemes. Performance variations across partitions are minimal, indicating that FedSEA-LLaMA is robust to different split configurations. This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
@@ Conclusion
In this work, we propose FedSEA-LLaMA, a novel federated split learning framework tailored for LLMs. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more base models and deploy the FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013,No.62388101, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
@@ Appendix A Noise Added to Hidden States
Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We use LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2. Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2. To assess the sensitivity to the noise scale when $Blocks_{A}$=3, we evaluate performance across a range of noise scales (from 0.1 to 0.4). As shown in Table 5, model performance decreases as the noise scale increases. This suggests that while moderate noise injection is tolerable, excessive noise harms representation quality.
Noise Scale F1a EM
0.1 75.68 34.10
0.15 71.21 26.82
0.2 59.08 16.47
0.3 55.79 7.45
0.4 57.68 6.09
Table 5: Performance under different noise scales when $Blocks_{A}$=3
@@ Appendix C Parallel Training
Appendix D Efficiency Comparison between FedSEA-LLaMA and FL-GLM
For the measurements of inference speed, inference memory usage, and per-token communication volume, we used the following prompt template:
{B_INST}{B_SYS}System: You are a helpful assistant. Always answer as helpfully as possible.{E_SYS}User: what can you do?{E_INST}
The total prompt length was 44 tokens. All inference experiments were conducted in FP16 precision with a maximum output length of 128 tokens. During inference, the client held one layer of model A and one layer of model C.
For the training memory usage experiments, we used the MultiRC dataset. Both model A and model C contained one layer. Unless otherwise specified, the default configuration was as follows: learning rate $1\times 10^{-5}$, LoRA rank =32=32, LoRA scaling factor =32=32, random seed =64=64, FP16 precision, and batch size =2=2. It is important to note that our training implementation actively clears unused variables in real time to minimize GPU memory usage. Therefore, the reported memory values correspond to the maximum observed usage within a one-minute monitoring window. All experiments were repeated three times with different random seeds (42, 64, and 128) to ensure the robustness and reproducibility of the results. The comparison results are shown in the Table 9.
Framework
Training Memory Usage
Inference Memory Usage
per-token communication
volume (Bytes)
Inference tokens
per second
FL-GLM 9810±0.00 6886±0.00 context length N × 10,335 0.75±0.03
FedSEA-LLaMA 5078±0.00 3674.00±0.00 8494 0.34±0.01
Table 9: Efficiency Comparison between FedSEA-LLaMA and FL-GLM
Appendix E Parameter Setting
@@
Ablation Study on Attention-Mask